P(4,0), m= -0.6
the problem asks me to find the equation in standard form using these points and the slope

Answer by stanbon(48564)

P(4,0), m= -0.6
the problem asks me to find the equation in standard form using these points and the slope
-----------
If y=mx+b, 0=(-6/10)(4)+b
0=-(24/10)+b
b=24/10
EQUATION:
y=(-6/10)x+24/10
Multiply thru by 10 to get:
10y=-6x+24
Put in standard form, as follows:
6x+10y=24
or, dividing thru by 2, you get:
3x+5y=12
Cheers,
Stan H.